the areas of the squares are at a rate of 3 to 7, if the smaller box has an area of 50 cm ^ 2. What is the area of the large square?
1
Expert's answer
2014-05-01T10:38:28-0400
Let the area of the large square is x the we get the equation
50/x=3/7 or 3x=350
Therefrom x=350/3 = 116.(6)=116+2/3 cm^2.